### v3

@@ 6 CONCLUSION
This study demonstrates that one-shot generative models, such as GPDM, significantly improve UAV classification metrics, especially in limited data scenarios where traditional deep generative models struggle. We rigorously evaluated the application of one-shot generative approaches in synthesizing RF signal transforms, establishing and validating an upper bound for the distributional distance between synthetic and real data. Our results consistently align with theoretical predictions across various scenarios, confirming the generation robustness in maintaining the integrity of data distributions. This validation ensures the synthetic data’s fidelity for reliable UAV identification. Furthermore, one-shot generative augmentation outperforms CGAN and CVAE in classifying UAVs and their flight mode with limited data and high signal complexity. Overall, our research offers a rigorous and effective approach for UAV identification in constrained data environments, paving the way for further exploration of modern generative models in cybersecurity.

### v4

@@ 7 Conclusion
This study demonstrates that one-shot generative models, such as GPDM, significantly improve UAV classification metrics, especially in limited data scenarios where traditional deep generative models struggle. We rigorously evaluated the application of one-shot generative approaches in synthesizing RF signal transforms, establishing and validating an upper bound for the distributional distance between synthetic and real data. Our results consistently align with theoretical predictions across various scenarios, confirming the generation robustness in distribution matching. This validation ensures the synthetic data’s fidelity for reliable UAV identification. Furthermore, one-shot generative augmentation outperforms CGAN and CVAE in classifying UAVs and their flight mode with limited data and high signal complexity. Overall, our research offers a rigorous and effective approach to improve UAV identification in limited RF environments.